You have given a function λ : R → R with the following properties (x ∈ R, n ∈ N):

λ(n) = 0 , λ(x + 1) = λ(x) , λ (n +1/2)=1

Find two functions p, q : R → R with q(x) not equal to 0 for all x such that λ(x) = q(x)(p(x) + 1). 


1
Expert's answer
2020-11-17T07:03:54-0500
"q(x)\not=0, x\in \R"

Then for "n\in \N"


"\lambda(n)=q(n)(p(n)+1)=0=>"

"=>p(n)+1=0=>p(n)=-1"

Let "q(x)=1,x\in \R."

Let "p(x)=|\sin(\pi x)|-1"

Then


"\lambda(x)=1\cdot(|\sin(\pi x)|-1+1)"

"\lambda(x)=|\sin(\pi x)|"

Check


"\lambda(x+1)=|\sin(\pi(x+1))|=|-\sin(\pi x)|="

"=|\sin(\pi x)|=\lambda(x), x\in\R"


"\lambda(n)=|\sin(\pi(n))|=0, n\in\N"

"\lambda(n+\dfrac{1}{2})=|\sin(\pi(n+\dfrac{1}{2} ))|=\sin(\dfrac{\pi}{2})=1, n\in\N"

"\lambda(x)=q(x)(p(x)+1)=1\cdot(|\sin(\pi x)|-1+1)"